# TIGeR: Tool-Integrated Geometric Reasoning in Vision-Language Models for Robotics

### v1

@@ Abstract
Vision-Language Models (VLMs) have shown remarkable capabilities in spatial reasoning, yet they remain fundamentally limited to qualitative precision and lack the computational precision required for real-world robotics. Current approaches fail to leverage metric cues from depth sensors and camera calibration, instead reducing geometric problems to pattern recognition tasks that cannot deliver the centimeter-level accuracy essential for robotic manipulation. We present TIGeR (Tool-Integrated Geometric Reasoning), a novel framework that transforms VLMs from perceptual estimators to geometric computers by enabling them to generate and execute precise geometric computations through external tools. Rather than attempting to internalize complex geometric operations within neural networks, TIGeR empowers models to recognize geometric reasoning requirements, synthesize appropriate computational code, and invoke specialized libraries for exact calculations. To support this paradigm, we introduce TIGeR-300K, a comprehensive tool-invocation–oriented dataset covering point transformations, pose estimation, trajectory generation, and spatial compatibility verification, complete with tool invocation sequences and intermediate computations. Through a two-stage training pipeline combining supervised fine-tuning (SFT) and reinforcement fine-tuning (RFT) with our proposed hierarchical reward design, TIGeR achieves SOTA performance on geometric reasoning benchmarks while demonstrating centimeter-level precision in real-world robotic manipulation tasks.
@@ II Related Work
Tool-Integrated Reasoning. Tool-Integrated Reasoning (TIR) enhances LLM and VLM reasoning ability, particularly for tasks requiring precise numerical computation, and is categorized into prompt-based, SFT-based, and RL-based methods. Prompt-based methods [13, 14, 15, 16, 17, 18, 19] use carefully designed prompts to invoke external tools without additional training. While easy to implement, they often exhibit low accuracy and unstable tool usage. SFT-based methods [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] fine-tune pre-trained models on TIR-specific data, improving performance but risking memorizing tool usage patterns from training data, which may hinder generalization to unseen scenarios. While current RL-based methods [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] promote generalization through outcome-based rewards, they overlook intermediate reasoning precision, essential for embodied tasks involving accurate numerical computation. We thus propose novel process-based reward functions tailored for TIR and introduce a new dataset to support training.
Spatial Understanding and Reasoning with VLMs. Spatial understanding focuses on object-centric properties (e.g., position, orientation) and inter-object relations (e.g., distance, direction), while spatial reasoning draws higher-level inferences over such information. Recent advances in VLMs enhance these two abilities via two paradigms: (1) Data-driven methods that fine-tune VLMs with pseudo-3D annotations [48, 8], real-world 3D datasets [49, 50], or simulated data [51, 52], using either the original image encoder, an added depth encoder [8, 11, 10] for 3D cues, or a geometry encoder [2, 1, 3] for structured scene and camera information. While these methods may enhance spatial understanding, they struggle with embodied tasks requiring precise numerical reasoning due to VLM’s hallucination and limited interpretability. (2) Tool-based methods [53, 54, 55] that integrate VLMs with vision foundation models to extract and reason spatial cues can achieve better performance but suffer from a limited training-free approach. In contrast, this work introduces a sequential SFT-RFT training pipeline with novel process reward functions, to enable spatial understanding and reasoning with accurate numerical computation.

### v2

@@ Abstract
Vision-Language Models (VLMs) have shown remarkable capabilities in spatial reasoning, yet they remain fundamentally limited to qualitative precision and lack the computational precision required for real-world robotics. Current approaches fail to leverage metric cues from depth sensors and camera calibration, instead reducing geometric problems to pattern recognition tasks that cannot deliver the centimeter-level accuracy essential for robotic manipulation. We present TIGeR (Tool-Integrated Geometric Reasoning), a novel framework that transforms VLMs from perceptual estimators to geometric computers by enabling them to generate and execute precise geometric computations through external tools. Rather than attempting to internalize complex geometric operations within neural networks, TIGeR empowers models to recognize geometric reasoning requirements, synthesize appropriate computational code, and invoke specialized libraries for exact calculations. To support this paradigm, we introduce TIGeR-300K, a comprehensive tool-invocation-oriented dataset covering point transformations, pose estimation, and spatial compatibility verification, complete with tool invocation sequences and intermediate computations. Through a two-stage training pipeline combining supervised fine-tuning (SFT) and reinforcement fine-tuning (RFT) with our proposed hierarchical reward design, TIGeR achieves SOTA performance on geometric reasoning benchmarks while demonstrating centimeter-level precision in real-world robotic manipulation tasks.
@@ II Related Work
Tool-Integrated Reasoning. Tool-Integrated Reasoning (TIR) enhances LLM and VLM reasoning ability, particularly for tasks requiring precise numerical computation, and is categorized into prompt-based, SFT-based, and RL-based methods. Prompt-based methods [13, 14, 15, 16, 17, 18, 19] use carefully designed prompts to invoke external tools without additional training. While easy to implement, they often exhibit low accuracy and unstable tool usage. SFT-based methods [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42] fine-tune pre-trained models on TIR-specific data, improving performance but risking memorizing tool usage patterns from training data, which may hinder generalization to unseen scenarios. While current RL-based methods [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56] promote generalization through outcome-based rewards, they overlook intermediate reasoning precision, essential for embodied tasks involving accurate numerical computation. We thus propose novel process-based reward functions tailored for TIR and introduce a new dataset to support training.
Spatial Understanding and Reasoning with VLMs. Spatial understanding focuses on object-centric properties (e.g., position, orientation) and inter-object relations (e.g., distance, direction), while spatial reasoning draws higher-level inferences over such information. Recent advances in VLMs enhance these two abilities via two paradigms: (1) Data-driven methods that fine-tune VLMs with pseudo-3D annotations [57, 8], real-world 3D datasets [58, 59], or simulated data [60, 61, 62], using either the original image encoder, an added depth encoder [8, 11, 10] for 3D cues, or a geometry encoder [2, 1, 3] for structured scene and camera information. While these methods may enhance spatial understanding, they struggle with embodied tasks requiring precise numerical reasoning due to VLM’s hallucination and limited interpretability. (2) Tool-based methods [63, 64, 65] that integrate VLMs with vision foundation models to extract and reason spatial cues can achieve better performance but suffer from a limited training-free approach. In contrast, this work introduces a sequential SFT-RFT training pipeline with novel process reward functions, to enable spatial reasoning with accurate numerical computation.